OpenStudy (anonymous):

solve the system: y=-x^2+5x+1 and y=2x+3. please tell me how you got your answer

OpenStudy (mertsj):

replace the y in the first equation with the value of y given by the second equation. 2x+3 = -x^2+5x+1 x^2 -3x +2 = 0 (x-2)(x-1)=0 x=2 or x = 1

hero (hero):

This is algebra 2 honors? You'd think it would be harder.

OpenStudy (anonymous):

we're having midterms and this is review for chapter 2 or 3 ok?

hero (hero):

You don't have to explain yourself

OpenStudy (anonymous):

how did you get x^2 isnt it supposed to be -x^2?

hero (hero):

I can explain it to you. vyew?

hero (hero):

?? Are you coming or not?

OpenStudy (anonymous):

no explain it here

hero (hero):

That site is an extension of OS

hero (hero):

Don't worry about the other stuff you see there. I will show you exactly how to do it.

OpenStudy (anonymous):

thats great, i dont want to go on that site though s

hero (hero):

Why not? What's wrong with it?

OpenStudy (anonymous):

are you gonna show me or not?

hero (hero):

Trust me, I'd rather show you on here, but the tools here aren't good enough.

hero (hero):

Someone already showed you but you still didn't understand.

OpenStudy (anonymous):

oh well ill figure it out on my own thanks for your time man

hero (hero):

Going on the other site will make sure you get it

hero (hero):

Oh well, your loss.

OpenStudy (anonymous):

no thanks man

OpenStudy (earthcitizen):

\[y=-x^2+5x+1=0\]

hero (hero):

The problem is, the OP can't see the problem being solved in real time. That's why he or she isn't understanding.

OpenStudy (earthcitizen):

OP ?

hero (hero):

Original Poster

OpenStudy (anonymous):

well then explain it

OpenStudy (earthcitizen):

can you find the roots of the equation ?

hero (hero):

I was but at a different location

OpenStudy (earthcitizen):

yh, Hero it's easy to explain

OpenStudy (anonymous):

yeah i got either (4,-1) or (-4,1)

OpenStudy (earthcitizen):

plug yhose values into the linear equation

OpenStudy (earthcitizen):

those are the points where the straight line and the curve meet

hero (hero):

Plug x back in to find y

hero (hero):

Obviously there are two solutions here

OpenStudy (anonymous):

3-1 doesnt get you 4...?

hero (hero):

Had to correct a slight mistake

hero (hero):

But anyways, there are two solutions. Plug x back in to find y

hero (hero):

No it isn't

OpenStudy (anonymous):

it would be -1 and -2

hero (hero):

It isn't the second step that's wrong. That's more like the fourth step, but okay

OpenStudy (anonymous):

yeah cause you fixed it but your solution is wrong

hero (hero):

Anyways, here's what you do: First set y = y: -x^2+5x+1 = 2x+3 Then move everything on the left side to the right side: x^2-3x+2 = 0 Then solve for x by factoring: x^2 - 3x+2 = 0 x^2-2x-1x +2 = 0 x(x-2)-1(x-2)=0 (x-2)(x-1) = 0 x = 2, x = 1

OpenStudy (anonymous):

wow i ended up doing the problem for you

hero (hero):

When we first started, you didn't even have a clue, so shush

OpenStudy (anonymous):

ma'am ma'am please

hero (hero):

Never complain when someone is trying to help YOU out

OpenStudy (anonymous):

oh and fyi bro you had like 3 mistakes in a 4 step equation so dont shush me

hero (hero):

That's because you didn't go to vyew like I asked you.

OpenStudy (anonymous):

hahahah someone even gave me a medal for good answer

OpenStudy (earthcitizen):

WTH

OpenStudy (anonymous):

alright ma'am, whatever floats your boat

hero (hero):

The bottom line is, I made mistakes trying to help you. It will never happen again.

OpenStudy (earthcitizen):

here